OpenStudy (anonymous):

Write the Equation of the line in standard form y=2/8x-3

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

same just subtract 2/8x both side standard form is Ax+By =C x and y should be on left side

OpenStudy (anonymous):

2/8x - y =3

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

i think

Nnesha (nnesha):

nope you have to subtract BOTH SIDE

Nnesha (nnesha):

wait no thats worng

OpenStudy (triciaal):



OpenStudy (anonymous):

when in standard form x cannot be negative

OpenStudy (anonymous):

so you flip the integers of everyone

OpenStudy (alexandervonhumboldt2):

add 3 to both side 3+y=2/8x subtract 2/8x from bth sides -2/8x+y=-3

Nnesha (nnesha):

nope you have to subtract both side by 2/8x bec its positive on left side

OpenStudy (triciaal):

no you multiply each side by -1 the signs may change not the integer meaning the number

OpenStudy (alexandervonhumboldt2):

2/8x-y=3 standart form

OpenStudy (alexandervonhumboldt2):

oh yah so

OpenStudy (alexandervonhumboldt2):

2/8x-y=3

OpenStudy (anonymous):

ty all